OpenStudy (anonymous):

how do i get y from 15y-8y=77

OpenStudy (anonymous):

(15-8)y = 77 so y = 77/(15-8)

OpenStudy (anonymous):

so y =7?

OpenStudy (anonymous):

or first subtract like terms in y then divide by coefficient of y

OpenStudy (anonymous):

no. 15-8 = 7 so y = 77/7

OpenStudy (anonymous):

oo ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw